OpenStudy (anonymous):

find b given that: A(4,10) and B(b,7) are 5 units apart. How can i solve it ?

OpenStudy (anonymous):

distance formula\[5=\sqrt(4-b)^2+(10-7)^2\] that sqrt should be over the whole right side

OpenStudy (anonymous):

then solve for b

OpenStudy (anonymous):

b=8 or b=0 there are 2 possibilities.

OpenStudy (anonymous):

So it would be 5 = 4-b + 9?

OpenStudy (anonymous):

could you show how you got your results?

OpenStudy (anonymous):

please cause i have 4 problems like this

OpenStudy (anonymous):

the (4-b) is still square and the right side should still be under the sqrt

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

\[5=\sqrt((4-b)^2+(10-7)^2)\] \[25=(4-b)^2+9\] subtract 9\[16=(4-b)^2\] take the squareroot, and remember that when taking the squareroot you have to use the negative and positive\[\pm4=4-b\]

OpenStudy (anonymous):

can you complete from there? or want me to finish?

OpenStudy (anonymous):

finish it please

OpenStudy (anonymous):

sorry! im just a better learning seeing it out solved

OpenStudy (anonymous):

so you have\[4=4-b \] subtract 4 and multiply by -1 to get rid of the negative \[0=b\] and you also have\[-4=4-b\] subtract 4\[-8=-b\] and multiply by -1 to get rid of the negative \[8=b\]

OpenStudy (anonymous):

I don't mind. Everyone learns differently and I'm happy to help.

OpenStudy (anonymous):

thank you but i have one more

OpenStudy (anonymous):

the question is

OpenStudy (anonymous):

A(b,b) is sqrt(18) units from its origin.

OpenStudy (anonymous):

find b

OpenStudy (anonymous):

the origin is at (0,0) so assuming the b is the same you have\[\sqrt18=\sqrt{(b-0)^2+(b-0)^2}\]

OpenStudy (anonymous):

so\[\sqrt{18}=\sqrt{b^2+b^2}\] \[\sqrt{18}=\sqrt{2b^2}\]square both sides \[18=2b^2\]divide by 2 \[9=b^2\]take the square \[\pm3=b\]root